Journalist Report – December 25th

Journalist Report 12-25-2017AD 7:33pm Sol 7
Author: John Sczepaniak MD

Today, was Christmas on Mars. In the morning the commander let us sleep an extra hour. Yay!!! We were priviliged to be visited by a three crew re-supply team and they brought us a wonderful Christmas meal. We had prime ribs, sugar coated sweet potatoes, butter coated biscuits, and ofcourse healthy vegitables. We were really stuffed. That was not all, we shared gifts from our respective countries and enjoyed French chocolates. For beverages with had apple cider (non-alcoholic). We shared our personal experiences and our plans once we return to Earth. It was a bit difficult for some because they would really like to be with their family back on that beautiful blue ball we call home. Alas, we will make due with the situation at hand; I would say we are doing well.

Overall it was nice to have a day off from extravehicular
activities. Going outside the habitat in a spacesuit is a bit physically draining. It requires everyone on the crew to prepare the suits, get participants dressed, and maintain communications. Not to mention the weight of the suit and the physical demands of wearing pressurized suits. The crew is now pretty accustomed to the decompression and repressurization. Our new space suit testing is coming along well and we hope to test it again tomorrow.

We wish everyone back on Earth a wonderful and bright Christmas. If you do not celebrate the holiday we wish you happy holidays and a beautiful New Year.

Greenhab Report – December 25th

Greenhab Report
David Murray
25 Dec, 2017

Environmental control:
Ambient
Ambient with door open from 9:30 to 17:18
Working Hour: 17:18
Inside temp at working hour: 16 C
Outside temp during working hours: 4.5 C
Inside temperature high: 32 C
Inside temperature low: 14 C
Inside humidity: 21%
Heating
Functioning nominally,
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
N/A
Changes to Crops
One spinach sprout
Morning Research Observations
N/A

Daily Water Usage for Crops
5 gallons
Time(s) of Watering Crops
17:18
Aquaponics
N/A
Narrative
One spinach seed has sprouted. All the plants look healthy. I’ll transplant the kale, cabbage, and spinach sprouts into larger pots over the next couple days.
Support/Supplies Needed
N/A

Journalist Report – December 25th

Journalist Report 12-25-2017AD 7:33pm Sol 7
Author: John Sczepaniak MD

Today, was Christmas on Mars. In the morning the commander let us sleep an extra hour. Yay!!! We were priviliged to be visited by a three crew re-supply team and they brought us a wonderful Christmas meal. We had prime ribs, sugar coated sweet potatoes, butter coated biscuits, and of course healthy vegetables. We were really stuffed. That was not all, we shared gifts from our respective countries and enjoyed French chocolates. For beverages with had apple cider (non-alcoholic). We shared our personal experiences and our plans once we return to Earth. It was a bit difficult for some because they would really like to be with their family back on that beautiful blue ball we call home. Alas, we will make due with the situation at hand; I would say we are doing well.

Overall it was nice to have a day off from extravehicular
activities. Going outside the habitat in a spacesuit is a bit physically draining. It requires everyone on the crew to prepare the suits, get participants dressed, and maintain communications. Not to mention the weight of the suit and the physical demands of wearing pressurized suits. The crew is now pretty accustomed to the decompression and repressurization. Our new space suit testing is coming along well and we hope to test it again tomorrow.

We wish everyone back on Earth a wonderful and bright Christmas. If you do not celebrate the holiday we wish you happy holidays and a beautiful New Year.

Sol Summary – December 25th

Sol 7

Merry Christmas!

David Murray

Mission Status: The crew is clean and well fed.

Sol Activity Summary: In the spirit of Christmas, Crew 185 rewarded ourselves with an extra hour of sleep and our first showers of the simulation. After a light breakfast, Arno and Ilaria left the hab to conduct an engineering check. Meanwhile, the rest of the crew prepared Christmas dinner, cleaned, and readied the hab to accommodate our guests. Shannon and her daughter arrived to share Christmas dinner with us. We had a first-rate feast of prime rib, candied yams, veggies, biscuits, soup, and baked potatoes. As much as we enjoyed the meal, we all miss our loved ones and wish that we were able to spend this holiday season with them. On behalf of Crew 185, MERRY CHRISTMAS!

Look Ahead Plan: The EVA for tomorrow focuses on GPS and drone control functionality in the X-1 spacesuit.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Healthy, clean, and well fed

EVA: X-1 tests: Drone Control & GPS

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, Sol Summary

Support Requested: None

Operations Report – December 25th

Operations Report

SOL: 7

Name of person filing report: A.Passeron

Non-nominal systems:

– Generator :
The generator worked properly during the night. It had a low battery failure at about 5am, which is normal since we charged the battery for only two hours yesterday. We keep going with the idea of charging the battery during the day to run the generator during the night. We started charging the battery at 8:40am. At 4:50pm, the charge of the battery was 75% while the SOC was 89%. We thus decided to connect the battery to the generator and it is working properly since then. At 7:20pm, the charge of the battery is still 12V.

Generator (hours run):
Generator turned off at roughtly 5:00am (failure)
Start charging battery at 8:40am
Battery charge is 75% at 4:50pm
Generator turned on at 4:50pm

Solar— SOC
@ 8:40am : 17%
@ 10:35am : 22%
@ 11:55am : 91%
@ 1:20pm : 87%
@ 1:45pm : 83%
@ 4:50pm : 89%
@ 7:20pm : 89%

Diesel – 76%

Propane – 50%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2.5 Gallons (1 container remaining)

Water (trailer) – 20 gallons

Water (static) – 280 gallons

Trailer to Static Pump used – No

Static to Loft Pump used – Yes

Water Meter: 50 gallons

Toilet tank emptied: Yes

ATVs Used: No
Oil Added?
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.

Deimos rover used: No
Number of hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: Assigned to director only.
Number of hours: 4.0
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Number of hours: 5.0
Beginning charge:
Ending charge:
Currently charging: Yes

Opportunity rover used: No
Number of hours: 4.3
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: No
Number of hours: 1.7
Beginning charge:
Ending charge:
Currently charging: Yes

HabCar used and why, where? No

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: None